A peach and a apple cost 70 cents. The peach is more expensive than the apple by 60 cents. What is the cost of each fruit?
Answer by Maths68(1474)

Let
Apple costs = x
Then
Peach costs =x+60
Apple cost + Peach Cost = Total Cost of two fruits
x+(x+60)=70
x+x+60=70
2x+60=70
2x=70-60
2x=10
Divide by 2 both sides
2x/2=10/2
x=5


Apple costs = 5 cents
Peach costs =x+60=5+60=65 cents